If I walk 20 miles north, then 15km east, then 10 km at 35 degree south of east, what is the magnitude of my displacement?


1
Expert's answer
2021-03-01T12:37:59-0500

Answer

Let initial point is "O"

If I walk 20 miles =32.2km north (let point=A) , then 15km east(point B) , then 10 km at 35 degree south of east(point C)

Now displacement OB from triangle OAB

"OB=\sqrt{OA^2+AB^2}"

="\sqrt{32.2^2+15^2}"

=35.52Km


Now final displacement

"OC=\sqrt{OB^2+BC^2}"

"=\sqrt{35.52^2+10^2}"

=36.9Km